please explain how to solve itHow much money will be in an account in 5 years if $10,000 is deposited now at an interest rate of 1% per month? Use three di
Answer #1

Formula for (F/P,i,n) = (1+i)^n

a.

per month interest rate = 1%

t = 5 * 12 = 60 months

Future worth = 10000 * (F/P,1%,60) = 10000 * (1+0.01)^60 = 10000 * 1.81669669 = 18166.97 ~ 18167 (Nearest Dollar)

b.

per month interest rate = 1%

Effective interest rate per quarter = (1+0.01)^3 - 1 = 1.0303 - 1 = 3.03%

t = 5 * 4 = 20 quarter

Future worth = 10000 * (F/P,3.03%,20) = 10000 * (1+0.0303)^20 = 10000 * 1.8166614 = 18166.61 ~ 18167 (Nearest Dollar)

c.

per month interest rate = 1%

Effective interest rate per quarter = (1+0.01)^12 - 1 = 1.126825 - 1 = 12.68%

t = 5 yrs

Future worth = 10000 * (F/P,12.68%,5) = 10000 * (1+0.1268)^5 = 10000 * 1.8166966 = 18166.96 ~ 18167 (Nearest Dollar)

Pls comment if you require further explanation